Personal Statement for Application to the Philosophy Graduate Program at the University of Oregon

George Moffitt

January 14, 2012

When I was in high school my older brother sent me a copy of Gödel, Escher, Bach: an Eternal Golden Braid. At the time I was focusing solely on mathematics, which made the book’s logic puzzles and hefty rational bias a tempting diversion. I had not ever read any philosophy, and was hard-pressed to even identify a single subject that the book could be about until long after I had finished it. The philosophy class I took at my Catholic high school was the cruel product of a pedagogical tyrant, a perspective that was more true the more clearly my later education brought it into focus. I had vowed at the end of high school to never take a philosophy class again, and spent my first year of college focusing almost exclusively on mathematics.

Linfield College required the usual liberal arts diversity credits, though, and mathematics satisfied none of them. I found myself signing up for an introductory philosophy class just to fill out my transcript, and discovered two important things: that introductory classes are extremely boring, and that even as I ignored the tepid class discussion I could not stop thinking about the material itself.

A friend convinced me to take a higher-level class with him, and I was immediately hooked on the culture of philosophical discourse. People who understood the text better and in a broader context than I were challenging me in a subject I quickly realized I adored. From there it was a slippery slope: I was a philosophy minor for only about a week before declaring my double major, and the more I studied the more philosophy eclipsed mathematics not just as an interest but as a lens for understanding my world. Ancient and Modern philosophy provided a perfect path of discovery for me, letting me study the philosophical positions of celebrated mathematicians as well as those of people who disagreed with them. I was still using philosophy to try to justify my mathematical bias, but this habit was gradually eroding as I found more problems that a mathematical view of the universe was ill-equipped to solve.

But I had begun too late, and while I had been exposed to post-modern thinking I left college still struggling with the problems of modernity. I graduated a skeptic, rereading Hume to try to make sense of experience and ethics. Considering myself not ready for graduate study in philosophy and no longer interested in mathematics, I got a job teaching swim lessons and wondered where to go next. I knew I needed to catch up on and really understand 20th-century philosophy, but while I tried working on Merleau-Ponty and Husserl I found little that addressed the basic philosophical understandings that were giving me trouble. I thought back to a colleague of mine at Linfield who had always been able to give me fresh perspective when I was stuck, patiently dissolving whatever assumptions were confining me. His work relied heavily on Heidegger, so I turned my attention to Being and Time. After two false starts, I sat down with a pad of paper and wrote a summary as I read. It took me a full year, but Heidegger’s refutation of Cartesian dualism and his novel phenomenological approach were the new understanding and direction I needed.

I have since been somewhat dedicated to studying Heidegger. His approach leaves no room for assumptions and admits (sometimes even embraces) his inaccuracies. This probably appeals to the mathematician in me, who keeps approximation and assumption to a minimum until the end of the problem before declaring them prominently. I think it is important for a philosopher to have Heidegger’s sense of discovery instead of merely the determination to prove a particular conclusion. Mathematics is a deductive system, and any conclusion you produce can only be a restating of your starting point. I would like to do more than rearrange the dust.

Surprisingly, the most important thing I’ve done these past years besides studying Heidegger has been my work as a swimming instructor. Having to teach something I thought I understood completely has been consistently challenging—I’ve learned that teaching requires more than an understanding of correct results, but a knowledge of every kind of wrong approach along the way as well as a Heideggerian relationship with truth that recognizes the need for individualized learning. Working with children has been similarly broadening in itself, showing me new ways of looking beyond my own perspective and giving me a real-time view of how changes in our society are affecting child development.

This has all renewed my resolve to become an educator in a more significant way. Additionally, I finally feel fundamentally ready for graduate study, to say nothing of desperately in need of a scholarly community in which to develop my own ideas. One thing that philosophy has instilled in me is a restlessness about believing the same thing for too long. I’ve been reading more diverse philosophers to try to gain a new perspective, but so far none has shaken, or even developed, the insights I’ve gained by reading Heidegger except for Heidegger himself. To that end, I am interested in performing my gruaduate studies in an environment where I can focus on Heidegger without being confined by him. I have always admired the ability great professors have to force me to take another look at philosophers I have dismissed, and while I do not wish to study analytic philosophy exclusively I also do not consider my mathematical background to be a waste. To that end I desire a place within a diverse department in which I can develop all the elements of my intellectual background with the phenomenological methods that have inspired me.

To that end, the University of Oregon’s commitment to philosophical diversity is particularly appealing to me. Its location in the Pacific Northwest is also important, since I grew up in Seattle and went to college in Oregon. Our region has significant environmental and social contexts in which I would like to continue grounding my education.

[Special Note: My letter of recommendation from Dr. Kaarina Beam at Linfield College will not be ready in time for the January 15th admission deadline, and potentially not even until early February. I appreciate your continued consideration pending the arrival of this crucial application material.]